OpenStudy (anonymous):

]Choose the equation of the line passing through the point (3, -2) and parallel to y = 3x - 6. y = 3x + 7 y = 3x - 11 y = -3x + 11 y = -3x - 12

OpenStudy (anonymous):

since it's parallel, you can eliminate two of those choices. and of those two, plug in the point (3, -2) and see which one works...

OpenStudy (anonymous):

how do you do that ?

OpenStudy (anonymous):

first, do you know which two to eliminate from consideration?

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

ok... in \(\large y=mx+b \), m is the slope of the line, and b is the y-intercept. in order for lines to be parallel, they must have the SAME slope..

OpenStudy (anonymous):

what is the slope of the line?

OpenStudy (anonymous):

hmmmm.... idk i'm all new to this. i'm sorry...

OpenStudy (anonymous):

your line is written in what's called slope-intercept form: \(\huge y=mx+b \) the number in front of the "x" is the slope of the line. that's what "m" is.

OpenStudy (anonymous):

this is your equation: y = 3x - 6 can you tell me what the slope is here?

OpenStudy (anonymous):

m=3, b=-6 where m is the slope slope The steepness of a line expressed as a ratio, using any two points on the line. and b is the y-intercept y-intercept The value of y at the point where a curve crosses the y-axis.

OpenStudy (anonymous):

ok... so the slope is 3, which of those choices have a slope of 3?

OpenStudy (anonymous):

all of them ?..

OpenStudy (anonymous):

be careful... there are only two with the SAME slope.

OpenStudy (anonymous):

i'm not really sure.

OpenStudy (anonymous):

what's the slope of the THIRD and LAST choice?

OpenStudy (anonymous):

11 and 12 ?

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

i was asking what is the SLOPE of the line in the third and last choices...

OpenStudy (anonymous):

i'm still talking about the number in front of the "x", the slope....

OpenStudy (anonymous):

3 and -3 ?

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

ok... the third and fourth choices have a slope of -3. it is these two that you can eliminate from consideration because the slope you have is 3. so only consider the first two choices because they both have a slope of 3.

OpenStudy (anonymous):

now, with the first two choices, plug in the point (3, -2) into each equation and see which one gives a true statement.

OpenStudy (anonymous):

plug in (3, -2) into each equation you get: for the first choice: y = 3x + 7 -2 = 3(3) + 7 continue to simplify... for the second choice: y = 3x - 11 -2 = 3(3) - 11 continue to simplify... which one gives a true statement in the end?

OpenStudy (anonymous):

i got it now! thank you so much for your time and help!!!!

OpenStudy (anonymous):

no problem...:) glad u stuck with it.... which one gave a correct statement?